OpenStudy (anonymous):

A principal of $3300 is invested at 6.5% interest, compounded annually. How much will the investment be worth after 14 years? Use the calculator provided and round your answer to the nearest dollar.

Nnesha (nnesha):

do you know the formula ?

OpenStudy (anonymous):

no i need help setting it up @Nnesha

OpenStudy (anonymous):

do i plug it in @RagingSquirrel

OpenStudy (ragingsquirrel):

Oh. r= annual interest rate

OpenStudy (ragingsquirrel):

Yes. Do you need help plugging the numbers in?

OpenStudy (anonymous):

yes it would be really helpful @RagingSquirrel

OpenStudy (phi):

I = Prt is simple interest. But you want compound interest

Nnesha (nnesha):

alright the formula is \[\large\rm A=P(1+ \frac{ r }{ n })^t\] p=principal (initial amount ) r=annual rate (we need in decimal so r/100) t= number of years A= amount after t years

Nnesha (nnesha):

compounded annually so n would be one

OpenStudy (anonymous):

i dont knw how to plug the numbers in @phi @Nnesha

Nnesha (nnesha):

P = what in ur question r= what ? divide by 100 bec we need in decimal t = number of years

OpenStudy (anonymous):

p=3300 r=6.5? t=14

OpenStudy (ragingsquirrel):

Convert 6.5% to a decimal. To do that, just move the decimal point two places. 0.065

Nnesha (nnesha):

right but r = 6.5 percent so we need to divide it by 100 and n =1 so plugin

OpenStudy (anonymous):

ok so what do I put for the t outside the parenthesis @Nnesha @RagingSquirrel

OpenStudy (anonymous):

the 14 @Nnesha

Nnesha (nnesha):

\[\large\rm A=P(1+ \frac{ r }{ n })^t\] p= 3300 t= 14 n=1 solve for A

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

ok im solving it now giv me a sec @Nnesha

Nnesha (nnesha):

sure take ur time!

OpenStudy (anonymous):

A= 7969.08 @Nnesha

OpenStudy (anonymous):

is that correct @Nnesha

Nnesha (nnesha):

yes right looks right `round your answer nearest dollar`

OpenStudy (anonymous):

thank you so much!!!!!! @Nnesha

Nnesha (nnesha):

np don't forget to round it!!